# Edge Graph Intelligence: Reciprocally Empowering Edge Networks with Graph Intelligence

### v1

@@ Abstract
Recent years have witnessed a thriving growth of computing facilities connected at the network edge, cultivating edge computing networks as a fundamental infrastructure for supporting miscellaneous intelligent services like personal voice assistance, video security surveillance, and autonomous driving vehicles. Meanwhile, Artificial Intelligence (AI) frontiers have extrapolated Machine Learning (ML) to the graph domain and promoted Graph Intelligence (GI), which unlocks unprecedented ability in processing, abstracting, and learning from massive data in graph structures. Given the inherent relation between graphs and networks, the interdiscipline of graph representation learning and edge networks, i.e., Edge GI or EGI, has revealed a novel interplay between them – GI models principally open a new door for modeling, understanding, and optimizing edge networks, and conversely, edge networks serve as physical support for training, deploying, and accelerating GI models. Driven by this delicate closed-loop, EGI can be widely recognized as a promising solution to fully unleash the potential of edge computing power and is garnering significant attention. Nevertheless, research on EGI yet remains nascent, and there is a soaring demand within both the communications and AI communities for a dedicated venue to share recent advancements. To this end, this paper promotes the concept of EGI, explores its scope and core principles, and conducts a comprehensive survey concerning recent research efforts on this emerging field and specifically, introduces and discusses: 1) fundamentals of edge computing and graph representation learning, 2) emerging techniques centering on the closed loop between graph intelligence and edge networks, i.e., “edge for GI” and “GI for edge”, and 3) open challenges and research opportunities of future EGI. By bridging the gap across communication, networking, and graph learning areas, we believe that this survey can garner increased attention, foster meaningful discussions, and inspire further research ideas in EGI.
@@ I Introduction
Figure 1: Illustration of the interplay between graph intelligence and edge networks, where graph intelligence can be applied as a data-driven tool to optimize edge networks, and conversely, edge networks perform as digital infrastructure to support graph intelligence deployment.
I-A Motivation and Benefits of Edge Graph Intelligence
@@
Given the remarkable success of Graph Intelligence (GI) and edge networks in their respective fields, the inherent connection between graphs and networks impels them to a confluence. As illustrated in Fig. 1, GI provides a vast zoo of empirical learning models (e.g., convolutional and recurrent GNNs, graph autoencoders) as well as various learning paradigms like Transfer Learning (TL) and Reinforcement Learning (RL), allowing advanced learning ability from graph data. Symmetrically, edge networks generally comprise of a rich set of platforms including mobile devices, robots, vehicles, and edge nodes, which host miscellaneous graph-based applications such as traffic forecasting and network resource management. Their bidirectional interaction, where GI enhances and optimizes edge networks and edge networks support and enable GI computation, draws a closed loop with mutual empowerment and nurtures a reciprocal interplay of their integration, namely “Edge Graph Intelligence” or “EGI” for brevity. More specifically, by fusing GI and edge computing, EGI provides three-fold reciprocal benefits in the following aspects.
  • Reciprocal performance enhancement: With the rapid proliferation of mobile and IoT devices, data generated at edge networks have skyrocketed in both quantity and modality (e.g., physical signals, digital audio, and visual content). As predicted by IDC [15], the billions of IoT devices in edge networks are expected to generate over 90 Zettabyte of data in 2025. This naturally provides a data nursery for modifying, training, and fine-tuning GI models with real-world data, thus boosting GI models toward higher-degree intelligence. Contrariwise, given the rich relational data collected at edge networks, GI enables modern graph analysis for understanding, diagnosing, and optimizing edge networks, which steers the enhancement of network performance such as robustness and Quality of Service (QoS).
  • Reciprocal capability expansion: As the last mile of the Internet, edge networks are continuously incubating pioneering user-centric scenarios surrounding terminal users, where many of them can be abstracted in graphs (e.g., wireless sensor networks, Internet of Vehicles). These scenarios act as a burgeoning force from the demand side to drive the development and deployment of GI, augmenting the application scope of GI models. Reversely, applying GI to edge networks thus unlocks their extended capability in securing edge networks from anomalies, developing new graph-based applications, and intelligently serving graph-related tasks. This is attributed to GI models’ innovative mechanism that combines graph embedding and convolution, which enables GI to deliver an unprecedented ability to learn and infer from graphs such as high-precision node identification and link prediction.
  • Reciprocal technology democratization: Technology democratization envisions making both GI and edge computing more usable, understandable, and friendly to a broader range of people, and has become a crucial agenda for social good. Towards that, GI and edge networks carry out mutual empowerment to improve each other. On the one hand, edge networks can democratize GI by rendering accessible and affordable intelligent computation closer to end users, allowing GI to be available and personalized in diverse scenarios. On the other hand, GI may popularize edge networks through prevailing GI-enabled applications at the edge. Specifically, given the computation-intensive and data-dependent nature of many GI models, edge computing exhibits its clear advantages against cloud computing by providing computing resources with lower delay and reduced bandwidth budget and is thus aligned well with GI. With the prevalence of GI surrounding end users, edge networks also gather attention and deployment for popularization.
I-B Scope and Rating of Edge Graph Intelligence
@@
While the term EGI is fresh to come, research and practices have begun early. Since the development of GCN in 2015 [16], GI has increasingly gained popularity in the AI community and ignited a wave of building GNNs over various real-world graphs. Meanwhile, edge networks and edge computing are also rapidly evolving and actively embracing AI from 2019, giving rise to the concept of edge AI or edge intelligence [8, 17, 18]. Currently, the interplay of EGI has attracted growing attention from both the industry and academia and propelled a plethora of innovative optimizations, techniques, and applications at the network edge, e.g., traffic flow forecasting [19, 20], location-based recommendation [21, 22], and vehicle trajectory prediction [23, 24]. As a substantial extension of edge AI, EGI sheds light on its fundamental questions - how deeply edge networks and AI techniques can be fused and how much potential their fusion can shine - and demonstrates its powerful capability through plentiful realistic applications.
Nonetheless, discussions on EGI are still narrowly limited to unilateral dimensions. Existing literature has either reviewed the graph learning landscape with limited discussion on their applications in edge networks [13, 9, 10], or focused particularly on applying GI techniques on some specific edge scenarios (e.g., traffic domain [25, 26, 27], power grids [28, 29, 30]), yet ignoring the big picture of general edge networks spectrum. Some recent literature [31, 32, 33] also reviews the progress of GI in the context of IoT and wireless networks but mainly focuses on GI applications in their discussed scopes and lacks a systematic taxonomy on the “Edge for GI” aspect, which is one of the fundamental pillars in EGI’s closed loop. Although the edge computing and communications community has extensively investigated edge intelligence systems, a majority of them [8, 34, 35, 36, 37] center on general AI computation systems or are dedicated to traditional DL workloads such as CNNs or RNNs, where GI models, which possess distinct capabilities and unique characteristics, are much less understood.
In this paper, we advocate that EGI should not be restricted to merely applying GI on edge data or running GI on edge platforms. Instead, GI and edge networks are blending a confluence and EGI should be treated as a whole to reflect the inherent interplay between GI and edge networks. This indicates that their bilateral empowerment desires a comprehensive exploration such that the degree of EGI can be identified and measured. Specifically, according to the fusion of GI and edge networks, we may rate EGI into six levels from their respective perspectives, as shown in Fig. 2.
  • Level 0: Given the graph data implicated in the edge network infrastructure, analytical models in Level 0 are unaware of the graph structures. The edge computing systems also process data in an graph-agnostic way. In other words, neither the model side nor the infrastructure side explicitly tackles “graphs”, and thus they are categorized to the initial level.
  • Level 1: Data collected from edge networks are modeled in graphs. Systems at Level 1 push one step further over Level 0 by endowing graph semantic to edge data (with general computing methods).
  • Level 2: Edge data in graph forms are processed with traditional graph computing algorithms such as PageRank and single-source shortest path algorithms. Systems at Level 2 outperform Level 1 by enabling the graph-oriented computing capability.
  • Level 3: Edge networks serve GI model inference with graph data, where the models may be trained on the cloud. Compared with lower levels, systems at Level 3 initiate AI to edge networks and embrace GI models.
  • Level 4: Edge networks perform GI model training with graph data. The key difference between Level 4 and Level 3 lies in the ability of learning edge-native GI models, e.g., fine-tuning model parameters with edge data.
The rating of EGI can be mainly divided into three intervals. The first interval covers from Level 0 to Level 2, where EGI is less related to AI and even processes non-graph data. The second interval comprises Level 3 and Level 4, where EGI incorporates GI models by either inference or training on edge networks. The third interval is exactly Level 5, which stands at the highest level because its GI and edge networks have profoundly blended as integration and can adapt to diverse scenarios on the fly. As EGI systems locates at higher levels, their fusions of GI and edge networks go deeper. As a result, the intelligence resources of GI and infrastructural resources of edge networks are progressively exploited for better EGI performance. Nonetheless, this may also come at the cost of additional development effort and system overhead. This conflict implies that there is no “silver bullet” in all cases. Instead, the panacea of EGI in practice should align with user demand, anticipating a joint consideration of specific application scenarios as well as available resources budgets.
I-C Summary and Contribution
@@
  • GI applications at Edge (Sec. IV): Typical application scenarios and use cases for applying GI in edge networks;
  • Edge Networks for GI (Sec. V): Paradigms of GI model computation, including model training and inference, for GI over edge networks;
  • GI for Edge Networks (Sec. VI): Practical GI-based methods for optimizing edge networks concerning their specific functionalities;
  • EGI ecosystems (Sec. VII): full-stack infrastructural support for high-performance EGI computation in terms of hardware, software, and benchmarks.
@@
Figure 3: Outline and conceptual relationships of EGI aspects discussed in this survey: Based on fundamental elements from GI (Sec. II) and edge computing networks (Sec. III), EGI ecosystems (Sec. VII) sustain all stakeholders in the closed loop of edge networks and graph intelligence. Edge Network for GI (Sec. V) reviews techniques for supporting edge computation of GI models and GI for Edge Network (Sec. VI) discusses GI-based optimizations on edge networks. Both of them serve as support for a rich set of EGI applications (Sec. IV).
@@
The rest of this paper is organized as follows: First, Sec. II and Sec. III briefly review the primers of graph intelligence and edge computing networks, respectively. Next, the subsequent sections introduce research efforts with respect to the four enablers: GI applications at Edge (Sec. IV), edge networks for GI (Sec. V), GI for edge networks (Sec. VI), and EGI ecosystems (Sec. VII). Finally, Sec. VIII discusses open challenges and future research opportunities of EGI and Sec. IX concludes. Table I lists the main abbreviations used in this survey.
II Primer on Graph Intelligence
@@
As one of the key flywheel actuating the loop within EGI , graph representation learning is devoted to the algorithm side and contributes enhanced ability in graph data processing. Before diving into EGI, this section compendiously introduces graph representation learning with respect to its basic concepts, general workflow, and representative models and learning paradigms. For a more comprehensive treatment of GI, concentrated reviews [38, 9, 13, 14] on graph representation learning are highly recommended. Table II lists the main notations used in this section.
II-A Basic Concepts
II-A1 Graphs
@@
Graphs are a way to organize data, and with graphs, one can succinctly characterize relationships across scattered data points. The input of GI models, i.e., GNNs, are graphs, which typically contain two types of data. One is the adjacency matrix or adjacency list, which interprets the graph topology, and the other is the feature vectors that describe vertices and edges’ actual properties. Formally, an input graph is denoted as $\mathcal{G}=\langle\mathcal{V},\mathcal{E}\rangle$, with vertices and links collected in $\mathcal{V}$ and $\mathcal{E}$, respectively.
II-A2 Vertices and Links
@@
Vertices or nodes in a graph can be items, objects, or entities, and are not necessarily homogeneous when constructing a graph. For instance, a location-based knowledge graph can represent its vertices as human users, IoT devices, scenic spots, and any other entities of various types within a specific district. Links are another essential component in graphs that characterizes the relationships between these items, objects, or entities. Note that to avoid misunderstanding, we exclusively use “links” to indicate the connection between vertices in an input graph while leaving “edge” for edge networks. A link can be defined with respect to the two (not necessarily unique) vertices associated with it. For $\mathcal{V}\in\mathcal{G}$ and $\mathcal{E}\in\mathcal{G}$, we denote their size, i.e., the number of vertices and links, as $V$ and $E$, respectively, and use $v$ and $e$ to index arbitrary vertex and link in them.
II-A3 Neighbors
@@
Neighbors are ego-networks centering on specific vertices within a graph. For a vertex $v$, its neighbors cover the vertices directly connected to $v$ and their adjoining links. Note that a vertex’s neighbors can be iteratively expanded by considering the neighbors of its neighbors. Formally, given $\mathcal{N}^{(k)}_{v}$ as vertex $v$’s $k$-hop neighbors, we have $\mathcal{N}^{(k+1)}_{v}=\{\mathcal{N}^{(1)}_{u}|\forall u\in\mathcal{N}^{(k)}_{v}\}$, where $\mathcal{N}^{(1)}_{v}$ indicates $v$’s one-hop direct neighbors.
II-A4 Representation Vectors, Features, and Embeddings
@@
Representation vectors are the numerical vectors associated with vertices and links, and are also referred to as encodings, representations, latent vectors, or high-level feature vectors depending on the context. In this section, we respectively denoted representation vectors by $h^{(l)}_{v}$ and $h^{(l)}_{e}$ at the $l$-th GNN layer. Upon input to the model, the initial representation vectors $h^{(0)}_{v}$ and $h^{(0)}_{e}$ are exactly the features attached to vertices and links, which quantify physical properties in specific applications. Extending the above knowledge graph example, the features of a vertex may include the users’ age and food preferences, and for a scenic spot, it can be location and popularity. After processing through model layers, the exported representation vectors are embeddings, a form of compressed feature representations of vertices, links, neighbors, or graphs. Embeddings can be viewed as the mappings of original data in latent spaces, which effectively reserve the semantics implicated in the input graph while can be used by downstream models for specific tasks (e.g., vertex classification and link prediction).
@@
II-A5 Model Output
The final output of GI models depends on the way of processing embeddings, i.e., the readout function. In general, the outcome of GI models can be grouped into three types: 1) Vertex-level output, where the result is vertex-wise predictions (e.g., classes, scores) for some dedicated vertices. 2) Link-level output, where the result is link-wise predictions for some dedicated links. 3) Graph-level output, where the results are the prediction of the whole graph (e.g., the operational status of a power grid).

### v2

@@ Abstract
Recent years have witnessed a thriving growth of computing facilities connected at the network edge, cultivating edge networks as a fundamental infrastructure for supporting miscellaneous intelligent services. Meanwhile, Artificial Intelligence (AI) frontiers have extrapolated to the graph domain and promoted Graph Intelligence (GI). Given the inherent relation between graphs and networks, the interdiscipline of graph learning and edge networks, i.e., Edge GI or EGI, has revealed a novel interplay between them – GI aids in optimizing edge networks, while edge networks facilitate GI model deployment. Driven by this delicate closed-loop, EGI is recognized as a promising solution to fully unleash the potential of edge computing power and is garnering growing attention. Nevertheless, research on EGI remains nascent, and there is a soaring demand within both the communications and AI communities for a dedicated venue to share recent advancements. To this end, this paper promotes the concept of EGI, explores its scope and core principles, and conducts a comprehensive survey concerning recent research efforts on this emerging field. Specifically, this paper introduces and discusses: 1) fundamentals of edge computing and graph learning, 2) emerging techniques centering on the closed loop between graph intelligence and edge networks, and 3) open challenges and research opportunities of future EGI. By bridging the gap across communication, networking, and graph learning areas, we believe that this survey can garner increased attention, foster meaningful discussions, and inspire further research ideas in EGI.
@@ I Introduction
Figure 1: Illustration of the interplay between GI and edge networks, where GI can be applied as a data-driven tool to optimize edge networks, and conversely, edge networks perform as digital infrastructure to support GI deployment.
@@
Given the remarkable success of Graph Intelligence (GI) and edge networks in their respective fields, the inherent connection between graphs and networks impels them to a confluence. As illustrated in Fig. 1, GI provides a vast zoo of empirical learning models (e.g., convolutional and recurrent GNNs, graph autoencoders) as well as various learning paradigms like Transfer Learning (TL) and Reinforcement Learning (RL), allowing advanced learning ability from graph data. Symmetrically, edge networks generally comprise of a rich set of platforms including mobile devices, robots, vehicles, and edge nodes, which host miscellaneous graph-based applications such as traffic forecasting and network resource management. Their bidirectional interaction, where GI enhances and optimizes edge networks and edge networks support and enable GI computation, draws a closed loop with mutual empowerment and nurtures a reciprocal interplay of their integration, namely “Edge Graph Intelligence” or “EGI” for brevity.
@@
While the term EGI is fresh to come, research and practices have begun early. Since the development of GCN in 2015 [15], GI has increasingly gained popularity in the AI community and ignited a wave of building GNNs over various real-world graphs. Meanwhile, edge networks and edge computing are also rapidly evolving and actively embracing AI since 2019, giving rise to the concept of edge AI or edge intelligence [8, 16, 17]. Currently, the interplay of EGI has attracted growing attention from both the industry and academia and propelled a plethora of innovative optimizations, techniques, and applications at the network edge, e.g., traffic flow forecasting [18, 19], location-based recommendation [20, 21], and vehicle trajectory prediction [22, 23]. As a substantial extension of edge AI, EGI sheds light on its fundamental questions - how deeply edge networks and AI techniques can be fused and how much potential their fusion can shine - and demonstrates its powerful capability through plentiful realistic applications.
@@
  • Edge applications of GI Models (Sec. V): Typical application scenarios and use cases for applying GI in edge networks;
  • Edge Networks for GI (Sec. VI): Paradigms of GI model computation, including model training and inference, for GI over edge networks;
  • GI for Edge Networks (Sec. VII): Practical GI-based methods for optimizing edge networks concerning their specific functionalities;
  • EGI ecosystems (Sec. VIII): full-stack infrastructural support for high-performance EGI computation in terms of hardware, software, and benchmarks.
@@
Figure 2: Outline and conceptual relationships of EGI aspects discussed in this survey: Based on fundamental elements of GI (Sec. III-A) and edge networks (Sec. III-B), EGI ecosystems (Sec. VIII) sustain all stakeholders in the closed loop of edge networks and graph intelligence. Edge Network for GI (Sec. VI) reviews techniques for supporting edge computation of GI models and GI for Edge Network (Sec. VII) discusses GI-based optimizations on edge networks. Both of them serve as support for a rich set of EGI applications (Sec. V).
@@
In summary, this paper makes the following key contributions.
  • A brief tutorial on the primer of EGI key components that explains basic and necessary concepts about GI and edge networks.
  • We submit an in-depth discussion on EGI advantages as well as an EGI rating taxonomy that can be used as a criterion to identify the advancement of EGI systems.
  • We conduct a comprehensive survey on various aspects of EGI, including edge applications of GI, edge networks for GI, GI for edge networks, and EGI ecosystems.
  • We discuss open challenges and research directions of EGI, acted as a reference for future EGI innovations.
@@
The rest of this paper is organized as follows: First, Sec. III-A and Sec. III-B briefly review the primers of graph intelligence and edge computing networks, respectively. Next, the subsequent sections introduce research efforts with respect to the four enablers: GI applications at Edge (Sec. V), edge networks for GI (Sec. VI), GI for edge networks (Sec. VII), and EGI ecosystems (Sec. VIII). Finally, Sec. IX discusses open challenges and future research opportunities of EGI and Sec. X concludes. Table I lists the main abbreviations used in this survey.
II Related Survey
While EGI intersects edge networks and GI, its investigation is still narrowly limited to unilateral dimensions.
On the GI side, a volume of literature has reviewed the general landscape of GL, providing insights on model taxonomy [24], capability boundaries [13], and practical tutorials [14]. As flourishing attention is attracted, publications relevant to GI topics explode. Following this trend, some researchers begin to deliver reviews concentrated on specific types of models, such as GAE [25, 26], STGNN [27, 28] and DGRL [29, 30]. Surveys on certain scenarios are also released. For instance, some researchers review the use of GI in traffic domain [31, 32, 33], discussing how traffic networks can be constructed as graphs and how traffic patterns can be resolved with GI; several surveys target power grids, where the electrical network are naturally graphs [34, 35, 36]. Nevertheless, these investigations either center on graph learning landscapes with limited discussion about their roles in edge networks [13, 9, 10], or focused particularly on applying GI techniques on some specific edge scenarios, yet ignoring the big picture of general edge networks spectrum. Some recent literature [37, 38, 39] also reviews the progress of GI in the context of IoT and wireless networks, aiming at exploiting GI for optimizing communication channels in IoT services. These works, however, mainly focus on GI applications in their discussed scopes and lack a systematic taxonomy on how GI can be computed in hierarchical edge networks, which is one of the fundamental pillars in EGI’s ecosystem.
On the edge AI side, surveys on the broader edge computing topics have been carried out for years, pertaining to visions and challenges [7], systems and tools [40], communication and IoT perspectives [41, 42, 43], etc. With the popularization of AI in edge networks, many researchers turned their attention to the potential and future directions of edge AI. For instance, Zhou et al. [8] discuss the motivation and advantages of edge intelligence along with a grading mechanism for edge intelligence by assessing the training and inference on edge platforms. Wang et al. [44] submit a comprehensive taxonomy of edge AI and analyze a set of open challenges toward future edge AI development. Although these surveys have extensively investigated edge intelligence systems, a majority of them [8, 44, 45, 46, 47] center on general AI computation or are dedicated to traditional DL workloads such as CNN or RNN. GI models, which possess distinct capabilities and unique computing characteristics, are much less understood in the edge AI context.
In summary, while the AI and edge computing communities have pushed their investigation to their respective frontiers, a thorough review of EGI, the combination of both lines, is still absent and desires actions.
III Primer on Edge Graph Intelligence
Before diving into various aspects of EGI, we briefly review the basic concepts and relevant techniques of GI and edge networks, respectively.
III-A Graph Intelligence
@@
As one of the key flywheel actuating the loop within EGI, graph representation learning is devoted to the algorithm side and contributes enhanced ability in graph data processing. Before diving into EGI, this section introduces graph representation learning with respect to its basic concepts, general workflow, and representative models and learning paradigms. For a more comprehensive treatment of GI, concentrated reviews [24, 9, 13, 14] on GL are highly recommended. Table II lists the main notations used in this section.
III-A1 Basic Concepts
@@
Graphs. Graphs are a way to organize data, and with graphs, one can succinctly characterize relationships across scattered data points. The input of GI models, i.e., GNNs, are graphs, which typically contain two types of data. One is the adjacency matrix or adjacency list, which interprets the graph topology, and the other is the feature vectors that describe vertices and edges’ actual properties. Formally, an input graph is denoted as $\mathcal{G}=\langle\mathcal{V},\mathcal{E}\rangle$, with vertices and links collected in $\mathcal{V}$ and $\mathcal{E}$, respectively.
@@
Vertices and Links. Vertices or nodes in a graph can be items, objects, or entities, and are not necessarily homogeneous when constructing a graph. For instance, a location-based knowledge graph can represent its vertices as human users, IoT devices, scenic spots, and any other entities of various types within a specific district. Links are another essential component in graphs that characterizes the relationships between these items, objects, or entities. Note that to avoid misunderstanding, we exclusively use “links” to indicate the connection between vertices in an input graph while leaving “edge” for edge networks. A link can be defined with respect to the two (not necessarily unique) vertices associated with it. For $\mathcal{V}\in\mathcal{G}$ and $\mathcal{E}\in\mathcal{G}$, we denote their size, i.e., the number of vertices and links, as $V$ and $E$, respectively, and use $v$ and $e$ to index arbitrary vertex and link in them.
@@
Neighbors. Neighbors are ego-networks centering on specific vertices within a graph. For a vertex $v$, its neighbors cover the vertices directly connected to $v$ and their adjoining links. Note that a vertex’s neighbors can be iteratively expanded by considering the neighbors of its neighbors. Formally, given $\mathcal{N}^{(k)}_{v}$ as vertex $v$’s $k$-hop neighbors, we have $\mathcal{N}^{(k+1)}_{v}=\{\mathcal{N}^{(1)}_{u}|\forall u\in\mathcal{N}^{(k)}_{v}\}$, where $\mathcal{N}^{(1)}_{v}$ indicates $v$’s one-hop direct neighbors.
@@
Representation Vectors, Features, and Embeddings. Representation vectors are the numerical vectors associated with vertices and links, and are also referred to as encodings, representations, latent vectors, or high-level feature vectors depending on the context. In this section, we respectively denoted representation vectors by $h^{(l)}_{v}$ and $h^{(l)}_{e}$ at the $l$-th GNN layer. Upon input to the model, the initial representation vectors $h^{(0)}_{v}$ and $h^{(0)}_{e}$ are exactly the features attached to vertices and links, which quantify physical properties in specific applications. Extending the above knowledge graph example, the features of a vertex may include the users’ age and food preferences, and for a scenic spot, it can be location and popularity. After processing through model layers, the exported representation vectors are embeddings, a form of compressed feature representations of vertices, links, neighbors, or graphs. Embeddings can be viewed as the mappings of original data in latent spaces, which effectively reserve the semantics implicated in the input graph while can be used by downstream models for specific tasks (e.g., vertex classification and link prediction).
@@
Model Output. The final output of GI models depends on the way of processing embeddings, i.e., the readout function. In general, the outcome of GI models can be grouped into three types: 1) Vertex-level output, where the result is vertex-wise predictions (e.g., classes, scores) for some dedicated vertices. 2) Link-level output, where the result is link-wise predictions for some dedicated links. 3) Graph-level output, where the results are the prediction of the whole graph (e.g., the operational status of a power grid).